# Sommerfeld Enhancement for Vector Dark Matter in the Fundamental Representation of $SU(2)_{L}$

### v2

@@ 2 The Model
Another important sector of the Lagrangian is the part of the potential which describes the interaction between the Higgs field and our massive vector doublet. Of course, the coupling constant that regulates the interactions between the Higgs field and the new vector field are free parameters. However, there is no reason to think that an important hierarchy arise among these couplings. Consequently, and for the sake of simplicity, we will take the assumption that all of them are of the same order $\lambda_{2}\approx\lambda_{3}\approx\lambda_{4}\approx\lambda$. In this way, $\lambda$ will be the only constant regulating the Higgs portal.
@@ 3.1 Basic Formalism
The so called “Sommerfeld enhancement” is a non-perturbative effect due to the resumation of initial states interactions. This effect may produce sensible consequences for the annihilation cross section in the non-relativistic regime [5, 19]. In the next few paragraphs we will review the formalism used to compute the Sommerfeld enhancement effect and the we will apply it to our model.
We start by considering the DM annihilation processes, which are described by a two-body state, in the non-relativistic regime. The non-relativistic Schwinger-Dyson equation for the Green function $\bra{0}T\phi_{b}(R^{\prime},\textbf{r}^{\prime})\phi_{a}^{\dagger}(R,\textbf{r})\ket{0}$ can be written as:
$\left[\left(i\partial_{R^{0}}+\frac{\nabla_{R}^{2}}{4M_{V}}+\frac{\nabla_{r}^{2}}{M_{V}}\right)-V(r)+i\frac{N}{2}\Gamma\delta^{3}(r)\right]$ $\bra{0}T\phi_{b}(R^{\prime},\textbf{r}^{\prime})\phi_{a}^{\dagger}(R,\textbf{r})\ket{0}$ (3)
$=i\delta^{4}(R-R^{\prime})\delta^{3}(\textbf{r}-\textbf{r}^{\prime})\delta_{ab}$
Here, $m$ represents the mass of Dark Matter, $\phi_{i}$ represents the DM field, $N$ is the number spin states for the two bodies system (for spin 1, we have $N=9$), $V(r)$ is electroweak matrices and $\Gamma$ is the so called annihilation matrix. The coordinate $=(R^{0},\textbf{R})$ R ) denotes the center-of-mass coordinate, and r represents the relative coordinate of the Dark Matter multiplets, defined as:
[…]
The imaginary part of the diagonal component of this Green function $\bra{0}T\phi_{b}\phi_{a}^{\dagger}\ket{0}$, is associated with the total annihilation cross section. We proceed to decompose the Green functions into individual partial wave modes.
$\bra{0}T\phi_{b}(R^{\prime},\textbf{r}^{\prime})\phi_{a}^{\dagger}(R,\textbf{r})\ket{0}=\int\dfrac{d^{4}P}{(2\pi)^{4}}e^{-iP\cdot(R-R^{\prime})}\sum_{l}\dfrac{2l+1}{4\pi}P_{l}(\cos\gamma)(-i)G_{ba}^{(E,l)}(r,r^{\prime})$ (5)
where $\gamma$ denotes the relative angle between r and $\textbf{r}^{\prime}$. The non-relativistic energy for a dark matter pair ($E$) is
$E=P^{0}-\dfrac{\textbf{P}^{2}}{4M_{V}}=\dfrac{M_{V}v^{2}}{4}$ (6)
where, $v$ represents the relative velocity between a pair of dark matter particles. By substituting (3) into (5), we obtain the following result:
$\left[-\dfrac{1}{M_{V}r}\dfrac{d^{2}}{dr^{2}}r-\dfrac{l(l+1)}{M_{V}r^{2}}-E+V(r)-i\Gamma\dfrac{N\delta(r)}{8\pi r^{2}}\right]G_{ba}^{(E,l)}(r,r^{\prime})=\dfrac{\delta(r-r^{\prime})}{r^{2}}\delta_{ba}$ (7)
In this analysis, we focus on the annihilation processes with orbital angular momentum $l=0$ (s-wave), which provide the leading-order contributions in the non-relativistic limit. Indeed, we verified that the $l=1$ contribution is at least one order of magnitude smaller than the leading s-wave. Consequently, in our calculations the total angular momentum $J$ is equal to the total spin angular momentum $S$. The leading-order potential terms exhibit spherical symmetry, resulting in the conservation of $J$ in annihilation processes. Furthermore, the conservation of total electric charge $Q$ is upheld. As a result, the Green function is expressed in the diagonal form for each eigenstate corresponding to $Q$ and $J$. To simplify the differential equation, we introduce $G_{ba}^{(E,l)}(r,r^{\prime})=\frac{g_{ba}(r,r^{\prime})}{rr^{\prime}}$.
$\left[-\dfrac{1}{M_{V}r}\dfrac{d^{2}}{dr^{2}}r-E+V(r)-i\Gamma\dfrac{N\delta(r)}{8\pi r^{2}}\right]g_{ba}^{(E,l=0)}(r,r^{\prime})=\delta(r-r^{\prime})\delta_{ba}$ (8)
The cross-section is determined by the imaginary part of the solution, and we derive the leading-order solution within the perturbative expansion of $\alpha_{2}$. The leading order solution can be written as
$g_{ba}(r,r^{\prime})|_{leadingorder}=\dfrac{iNm^{2}}{8\pi}\left[g_{>}(r^{\prime})\Gamma g_{>}^{T}(r)\right]_{ba}$ (9)
In the above expression, we introduce $g_{>ab}$ which is the solution of the following equation
$\left[-\dfrac{1}{M_{V}}\dfrac{d}{dr^{2}}-E+V(r)\right]g_{>ab}(r)=0$ (10)
Where $V(r)=V_{g-V}(r)+V_{h-V}(r)$ represents the electroweak potential, composed of $V_{g-V}(r)$ (interaction gauge bosons-VDM) and $V_{h-V}(r)$ (interaction Higgs boson-VDM) terms, $E$ denotes the energy of dark matter particles, $m$ stands for the mass of dark matter, and $g_{>ab}(r)$ denotes the wave function for $(r\gg 1)$, with the corresponding boundary conditions.
$g_{>ab}(r=0)=\delta_{ba}$ (11)
We impose different profiles depending on each potential. The short-range Yukawa potential diminishes as $r$ approaches infinity, causing the solution to exhibit an outgoing plane wave behavior. In the case of the long-range Coulomb potential, the asymptotic profile is determined by solving the Schrödinger equation in (10). In particular, we establish the following condition to govern the asymptotic behavior.
$g_{>ba}(r)\longrightarrow\ \left\{\begin{array}[]{lcc}d_{ba}(E_{a})e^{i\sqrt{M_{V}E_{b}}r}\hskip 28.45274ptshort-range~{}potential\\ \\ d_{ba}(E_{a})e^{i\sqrt{M_{V}E_{b}}r}e^{i\frac{M_{V}\alpha}{2\sqrt{M_{V}E_{b}}}\log{2\sqrt{M_{V}E_{b}}r}}\hskip 28.45274ptlarge-range~{}potential\\ \end{array}\right.$ (15)
Here, $E_{a}$ and $E_{b}$ denote the non-relativistic kinetic energy of $\phi_{a}$ and $\phi_{b}$, respectively. Given the constraints of numerical calculations, it is not feasible to let $r$ approach infinity. Therefore, we enforce these asymptotic behaviors at a sufficiently large value of $r=r_{f}$, ensuring that the numerical solutions become independent of $r_{f}$. The deviation from the asymptotic plane wave profile is quantified by $d_{ba}(E_{a})$, which is commonly referred to as the Sommerfeld Factor.
We numerically solve equation (10) to obtain the Sommerfeld factor as a function of the relative velocity for the initial state $v$. This effect becomes significant in the non-relativistic regime ($v\ll 1$), where the potential energy becomes comparable to the kinetic energy ($E\simeq\frac{M_{V}v^{2}}{4}$ ) in the Schrödinger equations. Conversely, the enhancement factors approach unity for ($v\approx 1$). Consequently, we can recover the perturbative cross-section, as obtained by the Born-Oppenheimer approximation, by replacing the Sommerfeld factor with $1$.
@@
3.2 Annihilation Cross Section
The optical theorem enables us to establish a connection between $Im(G_{aa}^{E,l=0}(r,r^{\prime}))$ and the s-wave spin-averaged, velocity-weighted annihilation cross-section $\braket{\sigma_{a}v}$. This connection results in the following formula.
[…]
In the equation above, there is no summation taken over $a$, and $\sigma_{a}v$ represents the annihilation cross section of $\phi_{a}\sim V^{i}V^{j}$. We introduce a prefactor denoted as $c_{a}$, defined as follows:
[…]
If we substitute the leading order solution given in (9), we obtain
$\braket{\sigma_{a}v}=c_{a}\dfrac{M_{V}^{2}}{k^{2}}\lim_{E\longrightarrow\frac{k^{2}}{M_{V}^{2}}}\left(E-\dfrac{k^{2}}{m^{2}}\right)^{2}[\mathcal{A}(k)~{}\Gamma~{}\mathcal{A}^{\dagger}(k)]_{aa}$ (21)
where we define
[…]
The asymptotic behavior $r\longrightarrow\infty$ is important to evaluate $\mathcal{A}_{ba}(k)$. Substituting (15) into (22), we obtain
[…]
$=d_{ba}(E)\left(-\dfrac{k}{M_{V}}\right)\left(E-\frac{k^{2}}{M_{V}}\right)^{-1}$ (24)
We replace into (21), we express the annihilation cross section as below
$\braket{\sigma_{a}v}\simeq c_{a}\sum_{b,a}d_{ab}(E)\Gamma_{bc}d_{ac}^{\ast}(E)$ (25)
Specifically, we are interested in the velocity-weighted annihilation cross section for the partial annihilation processes $V^{0}V^{0}\longrightarrow XX^{\prime}$, where $X$ represents any Standard Model particle.
$\braket{\sigma_{a}v}_{XX^{\prime}}=c\sum_{\alpha,\beta}\sum_{J,J_{z}}(\Gamma_{XX^{\prime}}^{J,J_{z}})_{\alpha\beta}~{}d_{3\alpha}(E)d_{3\beta}^{\ast}(E)$ (26)
Here, the index $3$ must be fixed; therefore, the Sommerfeld enhancement factor implies that we are considering the process $(V_{0}V_{0}^{\ast}\longrightarrow XX^{\prime})$, with $c$ is set to $2$ for initial states consisting of identical particles. We define $E$ as approximately $(E\simeq\frac{M_{V}v_{rel}^{2}}{4})$ representing the non-relativistic kinetic energy of the V-particles, $(\Gamma_{XX^{\prime}}^{J,J_{z}})_{\alpha\beta}$ is the Annihilation matrices, with $X$ is any standard model particle, and $d_{\alpha\beta}(E)$ with $(\alpha,\beta=1,2,3,4)$ as the Sommerfeld enhancement factor.
3.3 Non-Relativistic Limit
@@ Acknowledgments
This work was partially found by Fondecyt grant 1230110, ANID PIA/APOYO AFB220004, and ANID—The Millennium Science Initiative Program ICN2019_044. We are very thankful to Motoko Fujiwara and Patricio Escalona for discussions and suggestions that contributed significantly to improve the quality of this paper.
@@ A.1 Gauge-Dark Matter
We are required to calculate the electroweak matrix using the effective action derived from the Lagrangian (2), resulting in the following expression:
$S_{eff,1}=-\frac{1}{2}\int d^{4}xd^{4}y~{}J^{\mu\dagger}(y)\Delta_{0~{}\mu\nu}^{-1}(x,y)J^{\nu}(x)$ (49)
Now, we proceed to compute $S_{eff,1}$ as given in equation (49) in the non-relativistic limit, commencing with:
$\Delta_{0~{}\mu\nu}^{-1}(x,y)=-\int\dfrac{d^{4}q}{(2\pi)^{4}}e^{iq(x-y)}\begin{pmatrix}\frac{1}{q^{2}+i\epsilon}&0&0&0\\ 0&\frac{1}{q^{2}-M_{Z}^{2}+i\epsilon}&0&0\\ 0&0&\frac{1}{q^{2}-M_{W}^{2}+i\epsilon}&0\\ 0&0&0&\frac{1}{q^{2}-M_{W}^{2}+i\epsilon}\end{pmatrix}g_{\mu\nu}$ (50)
Next, we substitute the gauge currents and (50) into (49), resulting in:
$S_{eff,1}=\int\dfrac{d^{4}xd^{4}yd^{4}q}{2(2\pi)^{4}}e^{iq(x-y)}\left(\dfrac{J^{\mu}_{A}(x)J_{A~{}\mu}(y)}{q^{2}+i\epsilon}+\dfrac{J^{\mu}_{Z}(x)J_{Z~{}\mu}(y)}{q^{2}-M_{Z}^{2}+i\epsilon}+\dfrac{2J^{\mu}_{W^{+}}(x)J_{W^{-}~{}\mu}(y)}{q^{2}-M_{W}^{2}+i\epsilon}\right)$ (51)
Now, we need to determine the non-relativistic limit of (51). However, we have the relativistic gauge currents expressed as $J_{V}=(J_{V}^{0},\vec{J_{V}}),$ where $V$ represents any gauge boson. With this representation, each term of (51) can be expressed as:
$\dfrac{J^{\mu}_{V}(x)J_{V~{}\mu}(y)}{q^{2}-M_{V}^{2}+i\epsilon}=-\dfrac{J^{0}_{V}(x)J_{V}^{0}(y)}{\vec{q}^{2}+i\epsilon}+\dfrac{J^{0}_{V}(x)J_{V}^{0}(y)q^{0^{2}}}{(\vec{q}^{2}+i\epsilon)(q^{2}-M_{V}^{2}+i\epsilon)}+\dfrac{\vec{J}^{\mu}_{V}(x)\vec{J}_{V\mu}(y)}{q^{2}-M_{V}^{2}+i\epsilon}$ (52)
The third term in equation (52) involves magnetic interactions between currents. To account for non-relativistic currents, we require that $|\vec{J_{V}}|\ll J_{V}^{0}$, and as a result, this term can be neglected. The second term is associated with the retarded effects on the interactions between currents. Therefore, we must perform integration by parts on this term, leading to:
$\int$ $d^{4}xd^{4}yd^{4}q~{}e^{iq(x-y)}\dfrac{J_{V}^{0}(x)J_{V}^{0}(y)q^{0^{2}}}{(\vec{q}^{2}+M_{V}^{2})(q^{2}-M_{V}^{2}+i\epsilon)}$
=$\int$= ∫ $d^{4}xd^{4}yd^{4}q~{}e^{iq(x-y)}\dfrac{\dot{J}_{V}^{0}(x)\dot{J}_{V}^{0}(y)}{(\vec{q}^{2}+M_{V}^{2})(q^{2}-M_{V}^{2}+i\epsilon)}$ (53)
Using the condition $\mid\frac{1}{v_{i}}\frac{dv_{i}}{dt}\mid\ll M_{V_{0}},$ we can deduce that $|\dot{J}{V}^{0}|\ll J{V}^{0}$, and the retarded effects render this term negligible in the non-relativistic limit. Consequently, we need to integrate the first term.
$\int\dfrac{d^{4}q}{(2\pi)^{4}}\dfrac{e^{iq(x-y)}}{\vec{q}^{2}+M_{V}^{2}}=\delta(x^{0}-y^{0})\int\dfrac{d^{3}q}{(2\pi)^{3}}\dfrac{e^{iq(x-y)}}{\vec{q}^{2}+M_{V}^{2}}=\delta(x^{0}-y^{0})\dfrac{e^{-M_{V}|\vec{x}-\vec{y}|}}{4\pi|\vec{x}-\vec{y}|}$ (54)
We substitute equation (54) into equation (51), resulting in the following effective action in the non-relativistic limit:
@@
The first term in equation (A.1) accounts for the Coulomb potential among non-relativistic particles, whereas the remaining terms describe Yukawa potential associated with the exchange of $W$ and $Z$ bosons.
@@ B.1 Gauge-Dark Matter Interaction
The non-relativistic vector dark matter particles have the potential to annihilate into relativistic particles, specifically gauge bosons of $SU(2){L}$. In the framework of effective quantum field theory, these annihilations are depicted by the inclusion of an imaginary potential or an absorptive term in the action, which accounts for the disappearance of pairs of non-relativistic particles. These absorptive terms originate from $S_{eff,2}$ as outlined in equation (38). To provide an interpretation, we can visualize $S_{eff,2}$ in terms of Feynman diagrams. In this representation, the matrix $\Delta_{0}^{-1}$ characterizes the propagators of gauge fields, and each $\mathcal{Q}$ represents a vertex featuring two gauge fields and two vector dark matter fields, as described by Chowdhury et al. [4].
Figure 6: Feynman diagram of quartic terms
@@
$S_{eff,2}$ $=-iTr\log\left(\Delta_{0}^{-1}\mathcal{Q}\Delta_{0}^{-1}\mathcal{Q}\right)$
$=-4i\int d^{4}xd^{4}yd^{4}zd^{4}w\dfrac{d^{4}q}{(2\pi)^{4}}\dfrac{d^{4}k}{(2\pi)^{4}}\dfrac{e^{iq(x-y)}}{q^{2}+i\epsilon}\dfrac{e^{ik(w-z)}}{k^{2}+i\epsilon}Tr\left(\mathcal{Q}(y,z)\mathcal{Q}(w,x)\right)$ (63)
The factor of 4 corresponds to the trace of gauge fields’ indices, and our focus is on the imaginary part. To proceed, we consider the gauge bosons in the loop to be on-shell, as follows:
$\dfrac{1}{q^{2}+i\epsilon}\rightarrow-2\pi i\delta(q^{2})$ (64)
Therefore, we perform an integration over the time-like component of the 4-momentum, resulting in the transformation $d^{4}q\delta(q^{2})\rightarrow\frac{d^{3}q}{2|\vec{q}|}$ when the gauge particles are on-shell. The integration of 4-momentum is then expressed as:
$\int\dfrac{d^{4}q}{(2\pi)^{4}}\dfrac{d^{4}k}{(2\pi)^{4}}\dfrac{e^{iq(x-y)}}{q^{2}+i\epsilon}\dfrac{e^{ik(w-z)}}{k^{2}+i\epsilon}=-\dfrac{i}{8\pi}\delta^{4}(x-y)$ (65)
Then, replace into equation (B.1), we can obtain:
$S_{eff,2}=\frac{i}{2\pi}\int d^{4}xTr\begin{pmatrix}Q_{AA}(x)&Q_{AZ}(x)&Q_{AW}(x)&Q_{AW}^{\ast}(x)\\ Q_{AZ}(x)&Q_{ZZ}(x)&Q_{ZW}(x)&Q_{ZW}^{\ast}(x)\\ Q_{AW}^{\ast}(x)&Q_{ZW}^{\ast}(x)&Q_{WW}(x)&0\\ Q_{AW}^{\ast}(x)&Q_{ZW}^{\ast}(x)&0&Q_{WW}^{\ast}(x)\\ \end{pmatrix}^{2}$ (66)
@@
$S_{eff,2}$ $=\dfrac{i}{2\pi}\int~{}d^{4}x\left(Q_{AA}(x)^{2}+Q_{ZZ}(x)^{2}+2Q_{WW}(x)^{2}+2Q_{AZ}(x)^{2}\right)$
@@
$Q_{AA}^{2}(x)=\dfrac{e^{4}}{64}\left[(1-\tan\theta_{W})^{2}((V_{\nu}^{0})^{\ast}V^{0~{}\nu})^{2}+4(V_{\nu}^{-}V^{+~{}\nu})^{2}+4(1-\tan\theta_{W})(V_{\nu}^{0})^{\ast}V^{0~{}\nu}V_{\nu}^{-}V^{+~{}\nu}\right]$ (70)
$Q_{ZZ}^{2}(x)=\dfrac{g^{4}}{64}$ $[(\tan^{2}\theta_{W}\sin^{2}\theta_{W}-\sin\theta_{W}\cos\theta_{W})^{2}((V_{\nu}^{0})^{\ast}V^{0~{}\nu})^{2}+4\sin^{4}\theta_{W}(V_{\nu}^{-}V^{+~{}\nu})^{2}$
@@
$Q_{AZ}^{2}(x)$ $=\dfrac{e^{2}g^{2}}{64}[(\cos\theta_{W}-\sin\theta_{W})^{2}((V_{\nu}^{0})^{\ast}V^{0~{}\nu})^{2}+\dfrac{\cos^{2}(2\theta_{W})}{\cos^{2}\theta_{W}}{\cos\theta_{W}}(V_{\nu}^{-}V^{+~{}\nu})^{2}$
@@
with add angular momentum matrices
$S^{0,0}=-\dfrac{1}{\sqrt{3}}\begin{pmatrix}1&0&0\\ 0&1&0\\ 0&0&1\end{pmatrix}$ (74)
$S^{1,1}=\dfrac{1}{2}\begin{pmatrix}0&0&-1\\ 0&0&-i\\ 1&i&0\end{pmatrix}\hskip 28.45274ptS^{1,0}=\dfrac{1}{\sqrt{2}}\begin{pmatrix}0&i&0\\ -i&0&0\\ 0&0&0\end{pmatrix}\hskip 28.45274ptS^{1,-1}=\dfrac{1}{2}\begin{pmatrix}0&0&-1\\ 0&0&i\\ 1&-i&0\end{pmatrix}$ (75)
$S^{2,2}=\dfrac{1}{2}\begin{pmatrix}1&i&0\\ i&-1&0\\ 0&0&0\end{pmatrix}\hskip 28.45274ptS^{2,1}=\dfrac{1}{2}\begin{pmatrix}0&0&-1\\ 0&0&-i\\ -1&-i&0\end{pmatrix}\hskip 28.45274ptS^{2,0}=\dfrac{1}{\sqrt{6}}\begin{pmatrix}-1&0&0\\ 0&-1&0\\ 0&0&2\end{pmatrix}$
$S^{2,-1}=\dfrac{1}{2}\begin{pmatrix}0&0&1\\ 0&0&-i\\ 1&-i&0\end{pmatrix}\hskip 28.45274ptS^{2,-2}=\dfrac{1}{2}\begin{pmatrix}1&-i&0\\ -i&-1&0\\ 0&0&0\end{pmatrix}$ (76)
We are using polarization vector for photons as:
$\epsilon_{i}^{\pm}(k_{1})=\dfrac{1}{\sqrt{2}}\begin{pmatrix}\mp\cos\theta\\ -i\\ \pm\sin\theta\end{pmatrix}\hskip 28.45274pt\epsilon_{i}^{\pm}(k_{2})=\dfrac{1}{\sqrt{2}}\begin{pmatrix}\pm\cos\theta\\ -i\\ \mp\sin\theta\end{pmatrix}$ (77)

### v3

@@ 2 The Model
Another important sector of the Lagrangian is the part of the potential which describes the interaction between the Higgs field and our massive vector doublet. Of course, the coupling constant that regulates the interactions between the Higgs field and the new vector field are free parameters. However, there is no reason to think that an important hierarchy arise among these couplings. Consequently, and for the sake of simplicity, we will take the assumption that all of them are of the same order $\lambda_{2}\approx\lambda_{3}\approx\lambda_{4}\approx\lambda$. In this way, $\lambda$ will be the only constant regulating the Higgs portal. After the electroweak symmetry breaking, the physical mass of $V_{\mu}$ ($M_{V}$) receives contributions from the Proca term ($+m_{V}^{2}V_{\nu}^{\dagger}V^{\nu}$) but also from the interactions with the Higgs field.
@@ 3.1 Basic Formalism
The so called “Sommerfeld enhancement” is a non-perturbative effect due to the resumation of initial states interactions. This effect may produce sensible consequences for the annihilation cross section in the non-relativistic regime [5, 19]. In the next few paragraphs we will review the formalism used to compute the Sommerfeld enhancement effect [20] and the we will apply it to our model.
@@
3.1.1 The Two-Particle State and Schrödinger Equation
To describe the state of two DM-DM particles, which must be neutral in charge and CP-even, we consider the following combinations:
|$\ket{g}=\begin{pmatrix}v_{0}v_{0}\\ a_{0}a_{0}\\ v_{0}a_{0}\\ v_{+}v_{-}\end{pmatrix}$ (3)
The order of the components in the vector of equation (3) is arbitrary, and we can use different orderings for the pairs of these doublets.
Since this effect is non-relativistic, we must use the Schrödinger equation for two bodies, derived from the Schwinger-Dyson equation [5], which is:
$\left[-\dfrac{1}{M_{V}}\dfrac{d^{2}}{dr^{2}}-\frac{l(l+1)}{M_{V}r^{2}}-E+V(r)-\dfrac{iN\Gamma\delta(r)}{8\pi r^{2}}\right]g_{ab}(r)=0$ (4)
where index $a$ represents the initial state of the process and $b$ the final state and $r$ is the relative position. The wave function $g_{ab}$ describes the two particles in a doublet of $SU(2)_{L}$. As we have seen, the potential includes interactions with gauge and Higgs bosons, corresponding to Coulomb (long-range) and Yukawa (short-range) potentials. The kinetic energy of these particles is $E\approx\frac{M_{V}v^{2}}{4}$, where $v$ is the relative velocity between them. The $\Gamma$ term is the imaginary potential and corresponds to an absorptive part, arising in the low-energy limit of quartic interactions. Finally, $N$ is the number of spin states, given by $N=(2l+1)^{2}$, which in our case is 9.
We can consider polarized initial states in the general case, but the main contribution in the non-relativistic limit occurs in the s-wave, i.e., when $l=0$, where the total angular momentum is given solely by the spin contribution. Therefore, the Schrödinger equation is:
$\left[-\dfrac{1}{M_{V}}\dfrac{d^{2}}{dr^{2}}-E+V(r)-\dfrac{iN\Gamma\delta(r)}{8\pi r^{2}}\right]g_{ab}(r)=0$ (5)
To solve this equation (5), we need the boundary conditions given by [5]:
$g_{ab}(r\rightarrow 0)=\delta_{ab}$ (6)
$g_{ab}(r\rightarrow\infty)=d_{ab}(E)e^{i\sqrt{M_{V}E}r}e^{i\left(\frac{M_{V}\alpha}{2M_{V}E}\ln(2\sqrt{M_{V}E}r)\right)}$ (7)
where $\delta_{ab}$ is a Kronecker delta, $\alpha$ is the fine-structure constant, and $d_{ab}$ are the Sommerfeld enhancement factors. The asymptotic condition of (7) is valid for large distances encompassing the Coulomb and Yukawa potentials. The equation (5) can be solved numerically as it does not have an analytical solution. Once the solutions are determined, we use the optical theorem, which relates the imaginary part of the Green’s functions to the thermally averaged total annihilation cross-section [5, 7] of DM particles:
$\braket{\sigma v}_{XX^{\prime}}=c\sum_{a,b}\sum_{J,J_{z}}(\Gamma_{XX^{\prime}}^{J,J_{z}})_{ab}~{}d_{3a}(E)d_{3b}^{\ast}(E)$ (8)
where $c$ is a normalization constant (2 for $a=b$, 1 for $a\neq b$). The fixed index 3 is related to the definition of the vector $\vec{s}$ (14), to select dark matter annihilation $(V_{0}V_{0}\longrightarrow XX^{\prime})$ with $X$ and $X^{\prime}$ being particles of the Standard Model.
3.1.2 Non-Relativistic Limit
@@ Acknowledgments
This work was partially found by Fondecyt grant 1230110, grant PIIC 2021-II UTFSM, ANID BECA/DOCTORADO NACIONAL 21210616, ANID PIA/APOYO AFB220004, and ANID—The Millennium Science Initiative Program ICN2019_044. We are very thankful to Motoko Fujiwara and Patricio Escalona for discussions and suggestions that contributed significantly to improve the quality of this paper.
@@ A.1 Gauge-Dark Matter
In the context of non-relativistic dark matter, we need to determine the potentials that interact with these particles in a low-velocity regime. For this, we need to determine the real potential (cubic terms) of the Lagrangian (2) or effective action of $SU(2)_{L}$, which is the electroweak matrix, where the conserved currents interact with the gauge fields of the Standard Model, resulting in the following expression:
$S_{eff,1}\approx\int d^{4}x~{}\left[J_{\mu}^{W^{+}}(x)W^{+\mu}(x)+J_{\mu}^{W^{-}}(x)W^{-\mu}(x)+J_{\mu}^{Z}(x)Z^{\mu}(x)+J_{\mu}^{A}(x)A^{\mu}(x)\right]$ (31)
After integrating out the gauge bosons in the path integral, we obtain the effective action dependent on the currents. We need to calculate the conserved currents from the cubic terms of the Lagrangian (2); the details of the calculation are in  [4].
@@ B.1 Gauge-Dark Matter Interaction
Non-relativistic vector dark matter particles have the potential to annihilate into relativistic particles, specifically gauge bosons of $SU(2)_{L}$. In quantum mechanics (Schrödinger equation), these annihilations are represented by an imaginary potential or an absorptive term (quartic terms) in the action, which explains the disappearance of pairs of non-relativistic particles.
These absorptive terms originate from $S_{eff,2}$. To provide an interpretation, we can visualize $S_{eff,2}$ in terms of Feynman diagrams. In this representation, the matrix $\Delta$ characterizes the propagators of the gauge fields, and each $\mathcal{U}$ represents a vertex featuring two gauge fields and two vector dark matter fields, as described by Chowdhury et al. [4, 24].
@@
$S_{eff,2}$ $=-iTr\log\left(\Delta^{-1}\mathcal{U}\Delta^{-1}\mathcal{U}\right)$
$=\frac{i}{2\pi}\int d^{4}x~{}Tr\left[\begin{pmatrix}U_{AA}(x)&U_{AZ}(x)&U_{AW}(x)&U_{AW}^{\ast}(x)\\ U_{AZ}(x)&U_{ZZ}(x)&U_{ZW}(x)&U_{ZW}^{\ast}(x)\\ U_{AW}^{\ast}(x)&U_{ZW}^{\ast}(x)&U_{WW}(x)&0\\ U_{AW}^{\ast}(x)&U_{ZW}^{\ast}(x)&0&U_{WW}^{\ast}(x)\\ \end{pmatrix}\right]^{2}$ (40)
@@
$S_{eff,2}$ $=\dfrac{i}{2\pi}\int~{}d^{4}x\left(U_{AA}(x)^{2}+U_{ZZ}(x)^{2}+2U_{WW}(x)^{2}+2U_{AZ}(x)^{2}\right)$
@@
$U_{AA}^{2}(x)=\dfrac{e^{4}}{64}\left[(1-\tan\theta_{W})^{2}((V_{\nu}^{0})^{\ast}V^{0~{}\nu})^{2}+4(V_{\nu}^{-}V^{+~{}\nu})^{2}+4(1-\tan\theta_{W})(V_{\nu}^{0})^{\ast}V^{0~{}\nu}V_{\nu}^{-}V^{+~{}\nu}\right]$ (44)
$U_{ZZ}^{2}(x)=\dfrac{g^{4}}{64}$ $[(\tan^{2}\theta_{W}\sin^{2}\theta_{W}-\sin\theta_{W}\cos\theta_{W})^{2}((V_{\nu}^{0})^{\ast}V^{0~{}\nu})^{2}+4\sin^{4}\theta_{W}(V_{\nu}^{-}V^{+~{}\nu})^{2}$
@@
$U_{AZ}^{2}(x)$ $=\dfrac{e^{2}g^{2}}{64}[(\cos\theta_{W}-\sin\theta_{W})^{2}((V_{\nu}^{0})^{\ast}V^{0~{}\nu})^{2}+\dfrac{\cos^{2}(2\theta_{W})}{\cos^{2}\theta_{W}}{\cos\theta_{W}}(V_{\nu}^{-}V^{+~{}\nu})^{2}$